OpenStudy (anonymous):

solve for x in the two equations below using substitution: 30x-5y=150, 60x+5y=750

OpenStudy (anonymous):

\( 30x-5y=150 \implies y=6x-30\) Sub into the other equation \(60x+5(6x-30)=750 \) Solve for \(x\)

OpenStudy (amorfide):

there is no way i would do this by substitution sorry

OpenStudy (apoorvk):

30x-5y = 150, or, 30x-150 = 5y or, y = (30x-150)/5 or, y = 6x-30 ....(i) substitute the value of 'y' from 'i' above in the second equation provided, and solve x like any simple equation then. plug in the value of 'x' you get in 'i' above after that, to find out the corresponding 'y'.